### v3

@@ 4.5 Inference speed analysis on Tri-Former
In this section, we ran each model 10 times on the same dataset and averaged the number of pixels processed per second. The experimental results are shown in Figure 11, which demonstrate that the proposed Tri-Former enjoy clear advantages in both classification accuracy and inference speed. Compared with SSFTT, Tri-Former achieves 3$\times$ faster inference speed, while maintaining advantage in classification accuracy. On the Houston University dataset, Tri-Former improves classification accuracy by 10% compared to SpectralFormer while also achieving about a 20% increase in inference speed.
@@
4.6 Comparison between AGT and state-of-the-art methods
For evaluating the effectiveness of the proposed AGT strategy, we conducted a comprehensive comparison with several representative architecture tuning methods. The comparison methods involved in the evaluation are as follows: Fine-tune (adopted in LWNet Zhang et al. (2019) and Two-CNN Yang et al. (2017) for ConvNet methods), LORA Hu et al. (2022), Clip-adapter Gao et al. (2021) and LST Sung et al. (2022) (for transformer methods).
In this experiment, the basic model is first pretrained on the Salinas dataset acquired through the AVRIS sensor. In this process, 50 labeled pixels are randomly extracted from each class to build a pretraining set. Then, the model is fine-tuned with various strategies on three distinct target datasets captured by different sensors: AVRIS, ROSIS and ITRES-CASI. Within the tuning stage, each class is provided with 50 training samples. The experimental results are consolidated in Figure 12, where the x-axis denotes the employed tuning methods and the y-axis depicts the evaluation metric. According to the comparison, we can draw four conclusions:
  • Conventional fine tuning strategy tailored for ConvNets proves ineffective on the cross-sensor HSI data. As shown in Figure 12, the Fine-tune strategy not only fails to yield any improvement but also negatively affects the final results. We hypothesize that this phenomenon can be attributed to the fundamental differences between the information learned by ConvNets and transformer architectures. ConvNets learn a set of filter parameters from the training data. The process of feature extraction in ConvNets is static. Differing from this, transformer architectures learn relationships between pixels, capturing global dependencies in the data. The transformer is a data-driven model with a dynamic feature extraction process, making it more sensitive to differences between the pretraining dataset and the new target dataset. Unfortunately, there are obvious differences among cross-sensor HSI datasets, which consequently hinder the effectiveness of fine tuning.
@@ 4.6 Comparison between AGT and state-of-the-art methods
4.7 Experiments on homogeneous and heterogeneous HSI datasets
To further validate the effectiveness of the proposed method and investigate the influence of homogeneity and heterogeneity on tuning experiments, we conducted two sets of experiments on the Indian Pines, Salinas, Pavia University, and Pavia Center datasets. The first two datasets, Indian Pines and Salinas, are homogeneous and captured by AVIRIS, while the last two, Pavia University and Pavia Center, are homogeneous and captured by ROSIS. In the first set of experiments, the basic model was pretrained on the Salinas dataset and fine-tuned on the Indian Pines and Pavia University datasets under three configurations: 25, 50, and 75 training samples per class. The experimental results depicted in the radar charts in Figure 13 show that:
@@ 4.7 Experiments on homogeneous and heterogeneous HSI datasets
4.8 Using AGT as a bridge between RGB datasets and HSI datasets
RGB data is abundant and highly diverse. The barrier between RGB and HSI datasets is that they are cross-modality datasets. In this section, we combine HSI reconstruction method and the proposed AGT to build a bridge to cross this barrier.
More specifically, a pretrained HSI reconstruction model Li et al. (2022) is applied to convert 3-channel RGB image to 32-channel pseudo-HSIs. Then the pseudo-HSIs are used to pretrain the basic model. The RGB dataset used here is CIFAR [^7] https://www.cs.toronto.edu/~kriz/cifar.html. We randomly extract 25 samples from each category as tuning samples and take the rest as the testing samples. Experimental results are listed in Table 5.
Table 5: Experimental Results of Cross Sensors Transfer Learning from RGB Data
@@ 4.8 Using AGT as a bridge between RGB datasets and HSI datasets
4.9 Visual analysis on attention gate
Figure 14: Visual results of the gate attention selection process. (a) and (b) are two different samples.
In this section, we investigate whether the proposed attention gate can select relevant features from the basic model to enhance the features in the auxiliary branch. Specifically, following the inference process, we feed a large patch into the basic model and crop a smaller patch from its center to input into the auxiliary branch. The corresponding visual results are presented in Figure 14.
The visual results demonstrate that the attention gate functions passing relevant features selectively. If we directly connected the basic model and auxiliary branch, all features in the basic model be equally selected. Alternatively, if the proposed attention mechanism could not guide information flow from the basic model to the auxiliary branch, the selection map would appear random. However, Figure 14 shows that features from the basic model highly relevant to those in the auxiliary branch are selectively chosen. This observation underscores the effectiveness of the attention gate mechanism in fine tuning HSI classification methods.
5 Conclusion

### v4

@@ IV-E Inference speed analysis on Tri-Former
In this section, we ran each model 10 times on the same dataset and averaged the number of pixels processed per second. The experimental results are shown in Fig. 13, which demonstrate that the proposed Tri-Former enjoy clear advantages in both classification accuracy and inference speed. Compared with SSFTT, Tri-Former achieves 3$\times$ faster inference speed, while maintaining advantage in classification accuracy. On the Houston University dataset, Tri-Former improves classification accuracy by 10% compared to SpectralFormer while also achieving about a 20% increase in inference speed.
@@
Figure 15: Convergence curves of different architecture tuning strategies.
@@
IV-F Ablation study on Tri-Former
TABLE X: Ablation study
Components Performance (%)
Spectral module Spatial module 3D Conv OA AA K
- 97.26 97.31 96.83
98.06 97.97 97.76
97.26 97.05 96.84
98.41 97.87 98.17
To evaluate the function of each component in the proposed Tri-Former module, we conducted ablation study experiments on Indian Pines with 150 training samples each class. To ensure that the performance degradation is not due to the reduction in model parameters, the removed parts in the ablation experiment will be replaced with pointwise convolution or 3D convolution of equivalent size. The corresponding results are listed in Table X. From the experimental results, it can be concluded that each component of the design is necessary. Among them, the 3D convolution module has the greatest impact, which we attribute to its ability to simultaneously extract spatial-spectral features.
IV-G Comparison between AGT and state-of-the-art methods
For evaluating the effectiveness of the proposed AGT strategy, we conducted a comprehensive comparison with several representative architecture tuning methods. The comparison methods involved in the evaluation are as follows: Fine-tune (adopted in LWNet [15] and Two-CNN [16] for ConvNet methods), LORA [18], Clip-adapter [17] and LST [19] (for transformer methods).
In this experiment, the basic model is first pretrained on the Salinas dataset acquired through the AVRIS sensor. In this process, 50 labeled pixels are randomly extracted from each class to build a pretraining set. Then, the model is fine-tuned with various strategies on three distinct target datasets captured by different sensors: AVRIS, ROSIS and ITRES-CASI. Within the tuning stage, each class is provided with 50 training samples. The experimental results are consolidated in Fig. 14, where the x-axis denotes the employed tuning methods and the y-axis depicts the evaluation metric. According to the comparison, we can draw four conclusions:
  • Conventional fine tuning strategy tailored for ConvNets proves ineffective on the cross-sensor HSI data. As shown in Fig. 14, Fine-tune strategy not only fails to yield any improvement but also negatively affects the final results. We hypothesize this phenomenon can be attributed to the fundamental differences between the information learned by ConvNets and transformer architectures. ConvNets learn a set of filter parameters from training data. The process of feature extraction in ConvNets is static. Differing from this, transformer architectures learn relationships between pixels, capturing global dependencies in the data. The transformer is a data-driven model with a dynamic feature extraction process, making it more sensitive to differences between pretrain dataset and target dataset. Unfortunately, there are obvious differences among cross-sensor HSI datasets, which consequently hinder effectiveness of fine tuning.
@@ IV-G Comparison between AGT and state-of-the-art methods
  • Architecture tuning does have an impact on training effectiveness. To conduct a comprehensive analysis of different tuning strategies, we have plotted their respective convergence curves. As shown in Fig 15, compared with baseline, each tuning strategy exhibits a distinct convergence curve. Specifically, each tuning model achieves very high accuracy at beginning. The accuracy of training the base model (in blue line) gradually catches up with the accuracy curves of Fine-tune, Clip-adapter, and LORA. However, it still lags behind the convergence curve of AGT. We attribute this phenomenon mainly to the following reason. At the beginning of training, the baseline model starts with random initialization and lacks any feature extraction capability. Meanwhile, the tuned models have already learned feature extraction capabilities from heterogeneous data, allowing them to exhibit some effectiveness from the outset. During the training process, the baseline model benefits from the optimization on the new data, while tuned models may suffer from the conflicts gradually arising from the heterogeneous data. This dynamic change results in the baseline catching up with, and in some cases, surpassing the tuned models.
IV-H Experiments on homogeneous and heterogeneous HSI datasets
To further validate the effectiveness of the proposed method and investigate the influence of homogeneity and heterogeneity on tuning experiments, we conducted two sets of experiments on the Indian Pines, Salinas, Pavia University, and Pavia Center datasets. The first two datasets, Indian Pines and Salinas, are homogeneous and captured by AVIRIS, while the last two, Pavia University and Pavia Center, are homogeneous and captured by ROSIS. In the first set of experiments, the basic model was pretrained on the Salinas dataset and fine-tuned on the Indian Pines and Pavia University datasets under three configurations: 25, 50, and 75 training samples per class. The experimental results depicted in the radar charts in Fig. 16 show that:
@@ IV-H Experiments on homogeneous and heterogeneous HSI datasets
TABLE XI: Experimental Results of Cross Sensors Transfer Learning from RGB Data
@@
IV-I Using AGT as a bridge between RGB datasets and HSI datasets
RGB data is abundant and highly diverse. The barrier between RGB and HSI datasets is that they are cross-modality datasets. It is also important to note that both hyperspectral and RGB images are acquired by optical sensors, and thus, share similar spatial characteristics due to the common imaging mechanisms of light reflection, refraction, and diffraction. Moreover, most hyperspectral datasets inherently contain the RGB bands, and many hyperspectral sensors exhibit overlapping spectral ranges in these bands. These commonalities ensure that the spectral and spatial features of optical data are highly transferable. In contrast, SAR images, which are generated by active microwave radar systems, rely on scattering mechanisms that yield fundamentally different image characteristics. As a result, effective knowledge transfer between SAR and optical datasets maybe much more challenging.
In this section, we combine HSI reconstruction method and the proposed AGT to build a bridge to cross this barrier. More specifically, a pretrained HSI reconstruction model [46] is applied to convert 3-channel RGB image to 32-channel pseudo-HSIs. Then the pseudo-HSIs are used to pretrain the basic model. The RGB dataset used here is CIFAR [^9] https://www.cs.toronto.edu/~kriz/cifar.html. We randomly extract 25 samples from each category as tuning samples and take the rest as the testing samples. Experimental results are listed in Table XI.
@@ IV-I Using AGT as a bridge between RGB datasets and HSI datasets
IV-J Analysis on attention gate
Figure 17: Visual results of gate attention selection process. (a) and (b) show two different samples.
In this section, we investigate whether the proposed attention gate can select relevant features from the basic model to enhance the features in the auxiliary branch. Specifically, following the inference process, we feed a large patch into the basic model and crop a smaller patch from its center to input into the auxiliary branch. The corresponding visual results are presented in Fig. 17.
TABLE XII: Ablation study on the proposed attention gate
Methods Dataset Number of samples OA AA K
Addition Salinas to 50 92.71 96.71 91.68
AGT Indian Pines 50 94.83 97.62 94.07
The visual results demonstrate that the attention gate functions passing relevant features selectively. If we directly connected the basic model and auxiliary branch, all features in the basic model be equally selected. Alternatively, if the proposed attention mechanism could not guide information flow from the basic model to the auxiliary branch, the selection map would appear random. However, Fig. 17 shows that features from the basic model highly relevant to those in the auxiliary branch are selectively chosen. This observation underscores the effectiveness of the attention gate mechanism in fine tuning HSI classification methods. As shown in Table XII, if we replace the proposed attention gate architecture with a simple addition operation, the performance drops significantly, which also illustrates that the proposed attention gate is effective.
V Conclusion